# On estimation of weighted cumulative residual Tsallis entropy

### v1

@@ Abstract
Recently, weighted cumulative residual Tsallis entropy has been introduced in the literature as a generalization of weighted cumulative residual entropy. We study some new properties of weighted cumulative residual Tsallis entropy measure. Next, we propose some non-parametric estimators of this measure. Asymptotic properties of these estimators are discussed. Performance of these estimators are compared by mean squared error. Non-parametric estimators for weighted cumulative residual entropy measure are also discussed. Two uniformity tests are proposed based on an estimator of these two measures and power of the tests are compared with some popular tests. The tests perform reasonably well.
Keywords–asymptotic normality, consistency, order statistics, uniformity test
@@ 1 Introduction
As $\alpha\to 1$, it reduces to WCRE. Various properties of WCRTE measure along with its dynamic version and its potential as a risk measure for heavy tailed risk analysis was discussed in Chakraborty and Pradhan (2023). WCRTE is further studied by many authors, see, for example, Kattumannil et al. (2022, 2024) and Zuo and Yin (2025).
Estimations of entropy and related measures were considered in literature by many authors; see, for example, Vasicek (1976), Ebrahimi et al. (1994), Correa (1995), Noughabi (2010) and the references therein. Recently, research on weighted information measures gained quite a lot of attention see, for example, Balakrishnan et al. (2022), Chakraborty (2023), Hashempour et al. (2023), Khammar and Jahanshahi (2018) and the references therin. However, estimations of weighted measures have not been studied to the same extent. The main focus of this article is to study various non-parametric estimators of WCRTE measure. Chakraborty and Pradhan (2023) introduced one non-parametric estimator of WCRTE measure. Here we propose four estimators of WCRTE, study their asymptotic properties and compare their performances. Also, we briefly discuss non-parametric estimations of WCRE measure. We develop two uniformity tests based on an estimator of WCRTE and WCRE measures. The rest of the paper is organised as follows:
We study some important properties of WCRTE measure in Section 2. Non-parametric estimators of WCRTE and WCRE are introduced and their asymptotic properties are investigated in Section 3. A simulation study is performed to compare the performance of the estimators in Section 4. Uniformity tests are discussed in Section 5. Finally, some concluding remarks are made in Section 6.
@@ 2 Properties of WCRTE
Exponential $1-e^{-\lambda x};\;x>0,\;\lambda>0$ $\frac{\alpha+1}{\alpha\lambda^{2}}$
Rayleigh $1-e^{-\frac{x^{2}}{2\sigma^{2}}};\;x\geq 0,\;\sigma>0$ $\frac{\sigma^{2}}{\alpha}$
Pareto I $1-\left(\frac{k}{x}\right)^{\delta};\;x>k,\;\delta>0$ $\frac{\delta k^{2}}{(\delta-2)(\delta\alpha-2)}$, $\delta>(<)2,\alpha>(<)1$
Weibull $1-e^{-(\lambda x)^{p}},\;x>0,\;\lambda,\;p>0$ $\frac{\Gamma\left(\frac{2}{p}\right)\left(1-\frac{1}{\alpha^{2/p}}\right)}{p\lambda^{2}(\alpha-1)}$
@@ 3.1 Estimation of WCRTE
$\hat{\xi}_{\alpha}^{w}(X)$ $=$ $\frac{1}{2(\alpha-1)}\int_{0}^{+\infty}x\left(\bar{F}_{n}(x)-\bar{F}^{\alpha}_{n}(x)\right)dx$ (12)
@@ 4 Comparison of estimators
We conduct a simulation study to assess the performance of the proposed estimators. We compare the performance of these estimators by means of Bias and mean square error (MSE). For reference distributions, we consider exponential, Weibull and uniform distributions. We denote them by Exp($\lambda$), WE($\alpha,\beta$) and U($a,b$), respectively. The cdfs of these distributions are provided in Table 1. We generate 10000 random samples of size $n$ = 10, 20 and 30 from Exp(1), Exp(2), U(0,1) and WE(2,1) distributions and calculate bias and MSEs of these estimators. The results are provided in Tables 2-6.
Table 2: Bias and MSE of $\hat{\xi}_{\alpha}^{w}(X)$ and $\xi_{\alpha}^{w}L$ for $\alpha$ = 2.
$\hat{\xi}_{\alpha}^{w}(X)$ $\xi_{\alpha}^{w}L$
Distributions $n$ Bias MSE Bias MSE
10 -0.8259 0.9686 -0.7251 0.8848
Exp(1) 20 -0.7879 0.7751 -0.7400 0.7205
30 -0.7749 0.7143 -0.7427 0.6693
10 -0.2076 0.0616 -0.1814 0.0560
Exp(2) 20 -0.1964 0.0491 -0.1836 0.0451
30 -0.1940 0.0443 -0.1847 0.0416
10 -0.0086 0.00032 -0.0086 0.00043
U(0,1) 20 -0.0043 0.00014 0.0041 0.00017
30 -0.0029 0.00009 0.0028 0.00010
10 -0.0250 0.0078 0.0251 0.0102
WE(2,1) 20 -0.0120 0.0040 0.0116 0.0046
30 -0.0081 0.0027 0.0074 0.0030
Table 3: Bias and MSE of $\xi_{\alpha}^{w}V$, $\xi_{\alpha}^{w}E$ and $\xi_{\alpha}^{w}N$ for Exp(1) distribution with $\alpha$ = 2, $n$ = 10, 20 and 30 and for various $m$.
$\xi_{\alpha}^{w}V$ $\xi_{\alpha}^{w}E$ $\xi_{\alpha}^{w}N$
$n$ $m$ Bias MSE Bias MSE Bias MSE
10 1 -1.005 1.1386 -1.003 1.1301 -1.001 1.1314
@@
Table 4: Bias and MSE of $\xi_{\alpha}^{w}V$, $\xi_{\alpha}^{w}E$ and $\xi_{\alpha}^{w}N$ for Exp(2) distribution with $\alpha$ = 2, $n$ = 10, 20 and 30 and for various $m$.
$\xi_{\alpha}^{w}V$ $\xi_{\alpha}^{w}E$ $\xi_{\alpha}^{w}N$
$n$ $m$ Bias MSE Bias MSE Bias MSE
10 1 -0.2504 0.0713 -0.2497 0.0706 -0.2506 0.0707
@@
In Table 2, we present the bias and the MSEs of $\hat{\xi}^{w}_{\alpha}(X)$ and $\xi_{\alpha}^{w}L$ for these reference distributions. These two estimators do not require a window size. From Table 2, we observe that $\xi_{\alpha}^{w}L$ performs better than $\hat{\xi}^{w}_{\alpha}(X)$ for Exp(1) and Exp(2) distribution. But $\hat{\xi}^{w}_{\alpha}(X)$ performs better than $\xi_{\alpha}^{w}L$ for other two models. However, the difference is negligible for $n\geq 20$. Also, as sample size increases bias and MSE decrease.
The estimators $\xi_{\alpha}^{w}V$, $\xi_{\alpha}^{w}E$ and $\xi_{\alpha}^{w}N$ depend on the window size $m(<\frac{n}{2})$. We consider all possible values of $m$ for bias and MSE calculation and try to provide a guideline for the choice of $m$. The choice of $m$ will depend on the underlying distribution and the sample size $n$. We highlight the minimum values of MSEs of these estimators in Tables 3-6.
Table 5: Bias and MSE of $\xi_{\alpha}^{w}V$, $\xi_{\alpha}^{w}E$ and $\xi_{\alpha}^{w}N$ for U(0,1) distribution with $\alpha$ = 2, $n$ = 10, 20 and 30 and for various $m$.
$\xi_{\alpha}^{w}V$ $\xi_{\alpha}^{w}E$ $\xi_{\alpha}^{w}N$
$n$ $m$ Bias MSE Bias MSE Bias MSE
10 1 -0.0152 0.00050 -0.0145 0.00048 -0.0127 0.00043
@@
From Tables 3 and 4, it is observed that, for Exp(1) and Exp(2) distributions, $\xi_{\alpha}^{w}N$ performs better than other two estimators and $\xi_{\alpha}^{w}V$ performs the worst. All these estimators perform much better for Exp(2) distribution than Exp(1). For Exp(2) distribution, the difference between the MSEs of these estimators is negligible. The optimal choice of $m$ (for different $n$) for each estimator can be obtained from the respective tables. It is not possible to provide an exact formula for the choice of $m$ but approximate formulae may be provided from Tables 3-4, which will provide optimal or near optimal choices of $m$ for exponential distributions. Note from the tables that
  • (a)
  • (b)
    for $\xi_{\alpha}^{w}N$, choose $m=[\frac{n}{4}]+1$, for all $n$.
From Table 5 it is observed that, for U(0,1) distribution $\xi_{\alpha}^{w}N$ performs the best and $\xi_{\alpha}^{w}V$ performs the worst, although the performance of these estimators is similar. Performance of these estimators do not significantly vary for the choices of $m$ as well. Also, for U(0,1), these estimators perform better than all the other models considered. Here small values of $m$ will be appropriate.
Table 6: Bias and MSE of $\xi_{\alpha}^{w}V$, $\xi_{\alpha}^{w}E$ and $\xi_{\alpha}^{w}N$ for WE(2,1) distribution with $\alpha$ = 2, $n$ = 10, 20 and 30 and for various $m$.
$\xi_{\alpha}^{w}V$ $\xi_{\alpha}^{w}E$ $\xi_{\alpha}^{w}N$
$n$ $m$ Bias MSE Bias MSE Bias MSE
10 1 -0.0582 0.0081 -0.0562 0.0078 -0.0490 0.0072
@@
From Table 6, we observe that for WE(2,1) (Rayleigh) distribution, all estimators perform similarly, although the performance of $\xi_{\alpha}^{w}N$ may be considered to be the best whereas $\xi_{\alpha}^{w}V$ is the worst. However the difference in performance may be considered to be insignificant. Thus we prescribe to choose small $n$ for better result.
From the simulation study we find that $\xi_{\alpha}^{w}N$ performs the best, $\xi_{\alpha}^{w}V$ performs the worst and performance of $\xi_{\alpha}^{w}E$ stays in the middle for all the distributions considered. Also they perform better than $\hat{\xi}_{\alpha}^{w}(X)$ and $\xi_{\alpha}^{w}L$ in most of the cases. When sample size increases MSEs of all estimators decrease which is obvious. In general, $\xi_{\alpha}^{w}N$ has the lowest MSE among all other estimators. However, $\xi_{\alpha}^{w}N$ is more sensitive to the choice of $m$ than $\xi_{\alpha}^{w}V$ and $\xi_{\alpha}^{w}E$, whereas $\xi_{\alpha}^{w}V$ is the least sensitive to the choice of $m$. So, in order to work with $\xi_{\alpha}^{w}N$ and $\xi_{\alpha}^{w}E$, appropriate choice of $m$ is important for better results. Also, if the sample size is large then all estimators perform similarly. So, it is better to use $\xi_{\alpha}^{w}L$ for large $n$ because its asymptotic distribution is normal. This will be very useful in various statistical inference problems.

### v2

@@ Abstract
Recently, weighted cumulative residual Tsallis entropy has been introduced in the literature as a generalization of weighted cumulative residual entropy. We study some new properties of weighted cumulative residual Tsallis entropy measure. Next, we propose some non-parametric estimators of this measure. Asymptotic properties of these estimators are discussed. Performance of these estimators are compared by mean squared error. Non-parametric estimators for weighted cumulative residual entropy measure are also discussed. Estimator for weighted cumulative residual Tsallis entropy for progressive type-II censored data is proposed and its performance is investigated by Monte-Carlo simulations for various censoring schemes. Two uniformity tests for complete samples are proposed based on an estimator of these two measures and power of the tests are compared with some popular tests. The tests perform reasonably well. Uniformity test under progressively type-II censored data is also developed. Some real datasets are analyzed for illustration.
Keywords–asymptotic normality, consistency, order statistics, progressive type-II censoring, uniformity test
@@ 1 Introduction
As $\alpha\to 1$, it reduces to WCRE. Various properties of WCRTE measure along with its dynamic version and its potential as a risk measure for heavy tailed risk analysis was discussed in Chakraborty and Pradhan (2023). WCRTE is further studied by many authors, see, for example, Kattumannil et al. (2022, 2024), Zuo and Yin (2025) and the references therein.
Estimations of entropy and related measures were considered in literature by many authors; see, for example, Vasicek (1976), Ebrahimi et al. (1994), Correa (1995), Noughabi (2010), Al-Labadi et al. (2025) and the references therein. Recently, research on weighted information measures gained quite a lot of attention see, for example, Balakrishnan et al. (2022), Chakraborty (2023), Hashempour et al. (2023), Khammar and Jahanshahi (2018) and the references therin. However, estimations of weighted measures have not been studied to the same extent. The main focus of this article is to study various non-parametric estimators of WCRTE measure. Chakraborty and Pradhan (2023) introduced one non-parametric estimator of WCRTE measure. Here we propose four estimators of WCRTE, study their asymptotic properties and compare their performances. Also, we briefly discuss non-parametric estimations of WCRE measure. We introduce a non-parametric estimator for WCRTE under progressive type-II censored data. We develop two uniformity tests based on an estimator of WCRTE and WCRE measures for complete sample. Also, we develop an uniformity test for progressive type-II censored data as well. The rest of the paper is organised as follows:
We study some important properties of WCRTE measure in Section 2. Non-parametric estimators of WCRTE and WCRE are introduced and their asymptotic properties are investigated in Section 3. A simulation study is performed to compare the performance of the estimators in Section 4. A non-parametric estimator for WCRTE under progressively type-II censored data is discussed in Section 5. Uniformity tests are discussed in Section 6. Finally, some concluding remarks are made in Section 7.
@@ 2 Properties of WCRTE
Exponential $1-e^{-\lambda x};\;x>0,\;\lambda>0$ Exp($\lambda$) $\frac{\alpha+1}{\alpha\lambda^{2}}$
Rayleigh $1-e^{-\frac{x^{2}}{2\sigma^{2}}};\;x\geq 0,\;\sigma>0$ RA($\sigma$) $\frac{\sigma^{2}}{\alpha}$
Pareto I $1-\left(\frac{k}{x}\right)^{\delta};\;x>k,\;\delta>0$ PA($\delta,k$) $\frac{\delta k^{2}}{(\delta-2)(\delta\alpha-2)}$, $\delta>(<)2,\alpha>(<)1$
Weibull $1-e^{-(\lambda x)^{p}},\;x>0,\;\lambda,\;p>0$ WE($p,\lambda$) $\frac{\Gamma\left(\frac{2}{p}\right)\left(1-\frac{1}{\alpha^{2/p}}\right)}{p\lambda^{2}(\alpha-1)}$
@@ 3.1 Estimation of WCRTE
$\hat{\xi}_{\alpha}^{w}(X)$ $=$ $\frac{1}{(\alpha-1)}\int_{0}^{+\infty}x\left(\bar{F}_{n}(x)-\bar{F}^{\alpha}_{n}(x)\right)dx$ (12)
@@ 4 Comparison of estimators
We conduct a simulation study to assess the performance of the proposed estimators. We compare the performance of these estimators by means of Bias and mean square error (MSE). For reference distributions, we consider exponential, Weibull and uniform distributions. We denote them by Exp($\lambda$), WE($p,\lambda$) and U($a,b$), respectively. The cdfs of these distributions are provided in Table 1. We generate 10000 random samples of size $n$ = 10, 20 and 30 from Exp(1), Exp(2), U(0,1) and WE(2,1) distributions and calculate bias and MSEs of these estimators. The results are provided in Tables 2-5.
The estimators $\hat{\xi}^{w}_{\alpha}(X)$ and $\xi_{\alpha}^{w}L$ do not require a window size. However, $\xi_{\alpha}^{w}V$, $\xi_{\alpha}^{w}E$ and $\xi_{\alpha}^{w}N$ depend on the window size $m(<\frac{n}{2})$. We consider all possible values of $m$ for bias and MSE calculation and try to provide a guideline for the choice of $m$. The choice of $m$ will depend on the underlying distribution and the sample size $n$. We highlight the minimum values of MSEs of these estimators in Tables 2-5.
From Tables 2 and 3, it is observed that, for Exp(1) and Exp(2) distributions, $\xi_{\alpha}^{w}L$ performs the best for $n$ = 10 and $\xi_{\alpha}^{w}N$ performs the best for $n$ = 20 and 30. All these estimators perform much better for Exp(2) distribution than Exp(1). When the sample size increases the bias and the MSE of all the estimators decrease. The optimal choice of $m$ (for different $n$) for each estimator can be obtained from the respective tables. It is not possible to provide an exact formula for the choice of $m$ but approximate formulae may be provided from Tables 2-3, which will provide optimal or near optimal choices of $m$ for exponential distributions. Note from the tables that
  • (a)
  • (b)
    for $\xi_{\alpha}^{w}N$, choose $m=[\frac{n}{4}]+1$, for all $n$.
From Table 4 it is observed that, for U(0,1) distribution $\xi_{\alpha}^{w}N$ performs the best, although the performance of all five estimators is similar. Performance of $\xi_{\alpha}^{w}V$, $\xi_{\alpha}^{w}E$ and $\xi_{\alpha}^{w}N$ do not significantly vary for the choices of $m$ and small values of $m$ will be appropriate. Also, for U(0,1), these estimators perform better than all the other models considered. As there is no significant differnces bettwen the MSEs of these five estimators for U(0,1) model, any choice among these estimator will work. For simplicity, its best to work with $\hat{\xi}^{w}_{\alpha}(X)$ or $\xi_{\alpha}^{w}L$.
From Table 5, we observe that for WE(2,1) (Rayleigh) distribution, $\xi_{\alpha}^{w}N$ may be considered to be the best whereas $\xi_{\alpha}^{w}L$ is the worst. However the difference in performance may be considered to be insignificant. For $\xi_{\alpha}^{w}V$, $\xi_{\alpha}^{w}E$ and $\xi_{\alpha}^{w}N$, we prescribe to choose small $m$ for better result.
4.1 Data Analysis
We analyze a real dataset, given in Nelson (2005, page 105), to study the effectiveness of the proposed estimators. The following data describe failure times (in minutes) for an insulating fluid between two electrodes subject to a voltage of 34 kV.
0.19, 0.78, 0.96, 1.31, 2.78, 3.16, 4.15, 4.67, 4.85, 6.50,
7.35, 8.01, 8.27, 12.06, 31.75, 32.52, 33.91, 36.71, 72.89.
Here sample size is 19. The dataset has come from an exponential distribution with density
$f(x)=\lambda\exp(-\lambda x);\;x,\lambda>0.$
The maximum likelihood estimate (MLE) is $\hat{\lambda}=\frac{1}{\bar{X}}=0.0696$. We apply Kolmogorov-Smirnov test to check exponentiality of this dataset. The obtained p-value is 0.1672 which supports the exponentiality hypothesis. The theoretical value of WCRTE for exponential distribution is $\frac{\alpha+1}{\alpha\lambda^{2}}$.
We consider $\alpha$ = 2 for illustration. Then, the parametric estimate of WCRTE is $\frac{\alpha+1}{\alpha\hat{\lambda}^{2}}$ = 309.6512. We calculate non-parametric estimates for this dataset. For $m$-spacings estimators, we choose that $m$ which minimizes the absolute difference between the non-parametric estimates and the parametric estimate. The obtained values of the estimates are $\xi_{\alpha}^{w}(X)$ = 218.2337, $\xi_{\alpha}^{w}L$ = 232.5466, $\xi_{\alpha}^{w}V$ = 294.3883 for $m$ = 9, $\xi_{\alpha}^{w}E$ = 309.6309 for $m$ = 8 and $\xi_{\alpha}^{w}N$ = 304.8546 for $m$ = 5. So the estimators $\xi_{\alpha}^{w}E$ and $\xi_{\alpha}^{w}L$ perform well. Note that $m$-spacings estimators perform much better than $\xi_{\alpha}^{w}(X)$ and $\xi_{\alpha}^{w}L$.
For this dataset, Vasicek-type estimator performs best for maximum value of $m$, Ebrahimi-type estimator performs best for second largest choice of $m$ and $\xi_{\alpha}^{w}N$ performs best for $m$ = 5, which is the middle point. These values of $m$ are similar to the values found in Table 3 for Exp(1) distribution for $n$ = 20. They are in line with the choices provided in Section 4.
Table 2: Bias and MSE of the estimators for Exp(1) distribution with $\alpha$ = 2, $n$ = 10, 20 and 30 and for various $m$.
$\hat{\xi}_{\alpha}^{w}(X)$ $\xi_{\alpha}^{w}L$ $\xi_{\alpha}^{w}V$ $\xi_{\alpha}^{w}E$ $\xi_{\alpha}^{w}N$
$n$ Bias MSE Bias MSE $m$ Bias MSE Bias MSE Bias MSE
10 -0.8259 0.9686 -0.7251 0.8848 1 -1.005 1.1386 -1.003 1.1301 -1.001 1.1314
@@ 4.1 Data Analysis
Table 3: Bias and MSE of the estimators for Exp(2) distribution with $\alpha$ = 2, $n$ = 10, 20 and 30 and for various $m$.
$\hat{\xi}_{\alpha}^{w}(X)$ $\xi_{\alpha}^{w}L$ $\xi_{\alpha}^{w}V$ $\xi_{\alpha}^{w}E$ $\xi_{\alpha}^{w}N$
$n$ Bias MSE Bias MSE $m$ Bias MSE Bias MSE Bias MSE
10 -0.2076 0.0616 -0.1814 0.0560 1 -0.2504 0.0713 -0.2497 0.0706 -0.2506 0.0707
@@
Table 4: Bias and MSE of the estimators for U(0,1) distribution with $\alpha$ = 2, $n$ = 10, 20 and 30 and for various $m$.
$\hat{\xi}_{\alpha}^{w}(X)$ $\xi_{\alpha}^{w}L$ $\xi_{\alpha}^{w}V$ $\xi_{\alpha}^{w}E$ $\xi_{\alpha}^{w}N$
$n$ Bias MSE Bias MSE $m$ Bias MSE Bias MSE Bias MSE
10 -0.0086 0.00032 -0.0086 0.00043 1 -0.0152 0.00050 -0.0145 0.00048 -0.0127 0.00043
@@
Table 5: Bias and MSE of the estimators for WE(2,1) distribution with $\alpha$ = 2, $n$ = 10, 20 and 30 and for various $m$.
$\hat{\xi}_{\alpha}^{w}(X)$ $\xi_{\alpha}^{w}L$ $\xi_{\alpha}^{w}V$ $\xi_{\alpha}^{w}E$ $\xi_{\alpha}^{w}N$
$n$ Bias MSE Bias MSE $m$ Bias MSE Bias MSE Bias MSE
10 -0.0250 0.0078 0.0251 0.0102 1 -0.0582 0.0081 -0.0562 0.0078 -0.0490 0.0072
@@
5 Estimation of WCRTE for Progressively Type-II Censored Data
So far, we have discussed non-parametric estimation of WCRTE measure for complete sample. In this section, we propose an estimator for WCRTE measure for progressively type-II censored data. Censoring plays a crucial role in reliability and life-testing as it is not always possible to conduct experiments till all the units fail due to time and cost constraints. Type-I, type-II and hybrid censoring are commonly used in life-testing but these censoring schemes do not allow removal of items during the experiment. Cohen (1963) introduced progressive type-II (PC-II) censoring scheme which allows removal of units during the experiment. It is a generalization of the type-II censoring. A PC-II censoring can be described as follows:
Suppose $n$ identical units are put to a test, each having a common lifetime distribution with pdf $f$ and cdf $F$. Also, suppose a pre-fixed number $r$ of failures is allowed. Let $R_{1},\ldots,R_{r}$ be prefixed integers such that $R_{1}+\cdots+R_{r}=n-r$. When the first failure occurs, $R_{1}$ of the remaining n1n-1 surviving units are randomly removed from the test. Then $R_{2}$ of the remaining $n-R_{1}-2$ units are randomly removed at the time of second failure. Proceeding this way, at the time of the $r$th failure, all the remaining $R_{r}=n-r-\sum_{i=1}^{r-1}R_{i}$ units are removed from the test. The failure times $X_{1:r:n},\cdots,X_{r:r:n}$ are called progressive type-II censored order statistics. PC-II censoring reduces to the conventional type-II censoring when $R_{i}=0$ for $i=1,\ldots,r-1$, and $R_{r}=n-r$.
Let $F_{r:n}(x)$ denote the empirical distribution function of the PC-II censored data. Balakrishnan and Sandhu (1995) expressed $F_{r:n}(x)$ as
$F_{r:n}(x)=\begin{cases}0,\;\;\;\;\mbox{if}\;x<X_{1:r:n},\\ \alpha_{i:r:n},\;\;\;\mbox{if}\;X_{i:r:n}\leq x<X_{i+1:r:n},\;i=1,2,\cdots r-1,\\ \alpha_{r:r:n},\;\;\;\mbox{if}\;x\geq X_{r:r:n},\end{cases}$
where $\alpha_{i:r:n}=E(U_{i:r:n})$ and $U_{i:r:n}$ is the $i$th progressive type-II censored order statistic from U(0,1) distribution. Note that
$E(U_{i:r:n})=1-\prod_{k=r-i+1}^{r}\beta_{k},$
where $\beta_{i}=\frac{i+\sum_{k=r-i+1}^{r}R_{k}}{1+i+\sum_{k=r-i+1}^{r}R_{k}}$, $\beta_{k}=\beta_{1}$ if $k\leq 1$ and $\beta_{k}=\beta_{r}$ if $k\geq r$. For detailed discussions on PC-II censoring, readers may refer to Balakrishnan and Aggarwala (2000) and Balakrishnan and Cramer (2014).
Suppose $n$ items are put to a PC-II life-test and $X_{1:r:n},\cdots,X_{r:r:n}$ are the corresponding progressive type-II order statistics. We propose an estimator for WCRTE as
$\xi_{\alpha}^{w}P$ $=$ $\frac{1}{(\alpha-1)}\int_{0}^{X_{i:r:n}}x\left(\bar{F}_{r:n}(x)-\bar{F}^{\alpha}_{r:n}(x)\right)dx$ (23)
$=$ $\frac{1}{2(\alpha-1)}\sum_{i=1}^{r}\left(X^{2}_{i+1:r:n}-X^{2}_{i:r:n}\right)\left[\left(1-E(U_{i:r:n})\right)-\left(1-E(U_{i:r:n})\right)^{\alpha}\right].$
Proposition 5.1.
The estimator $\xi_{\alpha}^{w}P$ is consistent when $r\to n$, $n\to\infty$ and $\alpha>1$.
Proof.
When $r\to n$, the PC-II censored sample becomes complete sample and the required result follows from Lemma 6.2. ∎
Proposition 5.2.
For type-II censoring at $r$, the estimator reduces to
$\xi_{\alpha}^{w}P=\frac{1}{2(\alpha-1)}\sum_{i=1}^{r}\left(X^{2}_{i+1:r:n}-X^{2}_{i:r:n}\right)\left[\left(1-\frac{i}{n+1}\right)-\left(1-\frac{i}{n+1}\right)^{\alpha}\right].$
Proof.
The PC-II censoring becomes type-II censoring when $R_{i}=0$ for $i=1,\ldots,r-1$, and $R_{r}=n-r$. Now $\beta_{i}=\frac{i+n-r}{i+1+n-r}$ for all $i$ and $\prod_{k=r-i+1}^{r}\beta_{k}=\frac{n-i+1}{n+1}$. Substituting this in (23) we get the result. ∎
Table 6: Average Estimate and Variance of $\xi_{\alpha}^{w}P$ for U(0,1) and Exp(1) distributions.
U(0,1) Exp(1)
$n$ $r$ Schemes AE Var AE Var
20 10 (10,0910,0*9) 0.08236 0.00026 0.88841 0.61174
(09,100*9,10) 0.02271 0.00010 0.04970 0.00118
(5,08,55,0*8,5) 0.04188 0.00025 0.13144 0.00848
(1101*10) 0.06035 0.00050 0.29540 0.07601
15 (5,0145,0*14) 0.08470 0.00015 0.93225 0.44237
(014,50*14,5) 0.05593 0.00022 0.21008 0.01470
(07,5,070*7,5,0*7) 0.08487 0.00019 0.89215 0.50629
(3,013,23,0*13,2) 0.07328 0.00022 0.41973 0.06648
30 10 (20,0920,0*9) 0.08228 0.00027 0.87797 0.60185
(09,200*9,20) 0.00824 0.00002 0.01364 0.00009
(10,08,1010,0*8,10) 0.02180 0.00010 0.04666 0.00105
(2102*10) 0.03872 0.00038 0.12773 0.01706
20 (10,01910,0*19) 0.08501 0.00011 0.92785 0.34480
(019,100*19,10) 0.04610 0.00014 0.13597 0.00458
(5,018,55,0*18,5) 0.06410 0.00018 0.27141 0.01824
50 30 (20,02920,0*29) 0.08500 0.00007 0.91889 0.22099
(029,200*29,20) 0.03784 0.00008 0.09597 0.00142
(10,028,1010,0*28,10) 0.05880 0.00012 0.20955 0.00695
5.1 Simulation Study
To assess the performance of the proposed estimator we use Monte-Carlo simulation. We generate 5000 PC-II censored samples from U(0,1) and Exp(1) distributions and calculate average estimate (AE) and variance of $\xi_{\alpha}^{w}P$. We consider $(n,r)$ = (20,10), (20,15), (30,10), (30,20) and (50,30) and $\alpha$ = 2. The results are presented in Table 6 for various progressive type-II censoring schemes. In the table, the notation aba*b refers to ‘$a$ is repeated $b$ times’. For instance, (5,08,5)(5,0*8,5) refers to $(5,0,0,0,0,0,0,0,0,5)$. From the table, it is observed that when the sample size increases the variance of the estimator decreases for both the distributions. The overall performance of the estimator is good for both U(0,1) and Exp(1) distributions. As we have seen for complete samples, for PC-II censored data also, the estimator performs better for U(0,1) than Exp(1) distribution.
Chakraborty and Pradhan (2023) showed that WCRTE can be used as a risk (volatility) measure and they compared its performance with other popular risk measures such as standard deviation, right tail risk (Wang, 1998) using stock market (BSE SENSEX) data. Among all progressive censoring schemes, type-II censoring has the minimum volatility because it contains maximum information. From Table 6, we can see that for type-II censoring schemes the value of the AE of WCRTE is minimum and the corresponding variance of the estimator is also minimum for both the distributions. As the AE value increases, the variance of the estimator also increases. The PC-II censoring scheme where all the remaining items are removed after the first failure is called type-III censoring scheme (Pradhan and Kundu, 2009). Type-III censoring is opposite of type-II censoring and it has maximum volatility (minimum information) among all schemes. The AE values for type-III censoring are maximum, resulting in higher variance of the estimator $\xi_{\alpha}^{w}P$.